# Loop quantum effects on direct detection prediction in two-scalar dark matter scenario Karim Ghorbani

### v1

@@
We investigate the effect of quantum corrections on the elastic scattering cross section of dark matter off nucleus in two-scalar dark matter model. Among two extra singlet scalars in two-scalar model, the lighter one is stable and plays the role of dark matter candidate and the heavier one contributes in dark matter coannihilation processes in thermal history of the early universe. It is already known that the two-scalar model at tree level, in spite of the single-scalar dark matter model, can easily evade the bounds from direct detection (DD) experiments. The claim here is that taking into account the loop effects, in some regions of the parameter space, the DM-nucleon cross section becomes larger than the tree level contribution. Therefore, loop effects move the regions which were below the neutrino floor at tree level, up to the regions which are detectable by future DD experiments.
@@ 1 Introduction
There are noticeably two classes of models with dark matter candidates escaping the strong bounds from DD experiments. One scenario is that the scattering cross section of DM-nucleon tends to zero at tree level due to some symmetry breaking structure, for example, a pseudo-Goldstone boson as DM candidate in the complex scalar model with softly broken symmetry [7, 8, 9]. As such, in models with scale symmetry breaking, the DD cross section at tree level may be reduced significantly at some regions in the parameter space [10, 11]. The second avenue deals with models wherein the DM-nucleon scattring cross section is momentum or velocity suppressed, giving rise to DM candidates evading DD bounds easily. Among models of this type we can recall thermal DM candidates which interact with nucleons through pseudoscalar operators [12, 13, 14, 15, 16, 17, 18, 19].
@@ 2 Two-Scalar Model
$\phi_{1}=\varphi_{1}\cos\epsilon-\varphi_{2}\sin\epsilon\,,~{}~{}~{}\phi_{2}=\varphi_{1}\sin\epsilon+\varphi_{2}\cos\epsilon\,,$ (4)
@@
We take the scalar field $\phi_{2}$ with smaller mass $m_{2}$ as our DM candidate. We note that the interaction terms $\lambda_{1}\varphi_{1}^{4}$, $\lambda_{2}\varphi_{2}^{4}$, $\lambda_{12}\varphi_{1}^{2}\varphi_{2}^{2}$, $\lambda_{13}\varphi_{1}\varphi_{2}^{3}$ and $\lambda_{31}\varphi_{1}^{3}\varphi_{2}$, do not contribute to the DM annihilation cross section at tree level and also in the DM-nucleon cross section. However, in case we pick negative values for the couplings $\alpha_{1}$ and $\alpha_{2}$, to be assured of vacuum stability, we may set the couplings of the above five interaction terms at fix values other than zero. Since our plan in this research is to include the leading quantum corrections we will see that these five interaction terms do not play a role. Therefore in this model we are left with five independent free parameters: $m_{1},m_{2},\alpha_{1},\alpha_{2}$ and $\alpha_{12}$. The scalar mass difference defined as $\delta=m_{1}-m_{2}>0$ is used in our computation. In the following sections we may use the identity $m_{2}\equiv m_{\text{DM}}$.
As a requirement for the vacuum stability, the potential part of the Lagrangian should fullfil the bounded from below condition. The relevant formulas are found in [41] for a generic potential with two scalars:
@@
The parameters $a,b,c,d,e$ suited to our model read
$a=\frac{\lambda_{1}}{4}-\frac{\alpha_{1}^{2}}{4\lambda_{H}},~{}b=\lambda_{31}-\frac{\alpha_{1}\alpha_{12}}{\lambda_{H}},~{}c=\frac{\lambda_{12}}{2}-\frac{2\alpha_{12}^{2}+\alpha_{1}\alpha_{2}}{2\lambda_{H}},~{}d=\lambda_{13}-\frac{\alpha_{2}\alpha_{12}}{\lambda_{H}},~{}e=\frac{\lambda_{2}}{4}-\frac{\alpha_{2}^{2}}{4\lambda_{H}}.$ (8)
@@
where $\Gamma^{\text{SM}}_{higgs}$ is the Higgs decay width in the SM, $\Theta$ is the step fucntion and $m_{h}\sim 125$ GeV is the Higgs mass. The decay width of the Higgs in three different channels are provided by
$\Gamma(h\to\phi_{1}\phi_{1})=\frac{(\alpha_{1}\sin^{2}\epsilon+\alpha_{12}\sin\epsilon\cos\epsilon)^{2}v^{2}}{8\pi m_{h}}\sqrt{1-4m_{1}^{2}/m_{h}^{2}}\,,$ (10)
$\Gamma(h\to\phi_{2}\phi_{2})=\frac{(\alpha_{1}\cos^{2}\epsilon-\alpha_{12}\sin\epsilon\cos\epsilon)^{2}v^{2}}{8\pi m_{h}}\sqrt{1-4m_{2}^{2}/m_{h}^{2}}\,,$ (11)
$\Gamma(h\to\phi_{1}\phi_{2})=\frac{(\alpha_{1}\sin 2\epsilon+\alpha_{12}\cos 2\epsilon)^{2}v^{2}}{8\pi m_{h}^{3}}\sqrt{m_{h}^{2}-(m_{1}+m_{2})^{2}}\sqrt{m_{h}^{2}-(m_{1}-m_{2})^{2}}\,.$ (12)
The experimetal upper limit at the $95\%$ CL is found on the invisible Higgs decay, such that Br($h\to$ invisibles) $\lesssim 0.18$ [45]. This latter constraint becomes more effective for small mass of the scalars or large mixing angle.
@@ 3 Annihilation Cross Section and DM-nucleon cross section at Tree Level
The aim of this section is to reanalyze and update what is found within the present model in [37] from the calculations of the relic density and the DM-nucleon cross section at tree level. There are three channels through which the scalar DM can annihilate. 1) Through $s$-channel by Higgs particle as a mediator; with the SM fermions, gauge bosons and the Higgs in the final state. 2) Through $t$- and $u$-channel by one of the scalars as mediator; with a pair of the Higgs particles in the final state. 3) Annihilation into a pair of Higgs particles via a contact interation. In the left panel of Fig. 1, the corresponding Feynman diagrams for the DM annihilation are shown. We have not shown diagrams with more than two particles in the final state (i.e., $\phi_{2}\phi_{2}\to hhh$). To get co-annihilation diagrams one can simply replace one $\phi_{2}$ by one $\phi_{1}$ in the initial states. In the right panel of Fig. 1 the diagram for the DM-nucleon elastic scattering is depicted.
@@
Let us assume for a moment that $\alpha_{12}=0$ (thus $\epsilon=0$). This will turn the model back into its simplest form, only one singlet scalar in the model. In this simple case the annihilation cross section contains terms each of them proportional to $\alpha_{2}^{2}$, $\alpha_{2}^{3}$ or $\alpha_{2}^{4}$. Thus the DM relic density is inversly proportional to this coupling. In Appendix A, the annihilation cross sections are given for the full model. Now looking at the relevant vertex one can easily see that the DM-nucleon cross section depends on the same coupling, $\alpha_{2}$, see Appendix B. Since the DM-nucleon cross section and DM relic density dependency on the couplings go in opposite way, when varying the coupling $\alpha_{2}$, as it is shown in [37], it is not possible to respect both the DD bounds and the observed relic density at the same time (except for a resonance region where $m_{2}\sim m_{h}/2$).
@@
Now set $\alpha_{12}\neq 0$, and we get back to our present model. This brings in new interactions which give pivotal contribution to the DM annihilation cross section. We first identify three couplings related to the scalar-Higgs interactions. The strength of the interaction vertex $\phi_{2}\phi_{2}h$ is equal to $v\kappa_{22}$ with $\kappa_{22}=\alpha_{1}\cos^{2}\epsilon+\alpha_{2}\sin^{2}\epsilon-2\alpha_{12}\cos\epsilon\sin\epsilon$, the strength of the interaction vertex $\phi_{1}\phi_{2}h$ is equal to $v\kappa_{12}$ with $\kappa_{12}=\alpha_{12}\cos 2\epsilon+\cos\epsilon\sin\epsilon(\alpha_{1}-\alpha_{2})$, and finally the strength of the interaction vertex $\phi_{1}\phi_{1}h$ is $v\kappa_{11}$ with $\kappa_{11}=\alpha_{2}\cos^{2}\epsilon+\alpha_{1}\sin^{2}\epsilon+2\alpha_{12}\cos\epsilon\sin\epsilon$. Now the DM-nucleon cross section at tree level is only dependent on the coupling $\kappa_{22}$ while the DM annihilation cross section depends on the three couplings $\kappa_{11}$, $\kappa_{22}$ and $\kappa_{12}$. As a result, new viable regions in the parameter space might show up. The reason hinges on the fact that by incorporating these new contributions it becomes feasible to get small DM-nucleon cross section (small $\kappa_{22}$) respecting the current DD bounds, and in order to get the DM relic density right one can regulate the couplings $\kappa_{12}$ and $\kappa_{11}$ accordingly.
@@
$0\leq\alpha_{1}\leq 1,~{}~{}~{}0\leq\alpha_{12}\leq 2,~{}~{}~{}1\leq\delta\leq 100~{}(\text{GeV}),~{}~{}~{}30\leq m_{2}\leq 2000~{}(\text{GeV})\,.$ (13)
We present our numerical results for the DM-nucleon cross section in terms of the DM mass for $\alpha_{2}=0.25$, $\alpha_{2}=1$ and $\alpha_{2}=2$ in Figs. 2-4, while all the constraints mentioned above are applied. In these plots some new features appear when DM mass exceeds $\sim 125$ GeV. These were absent in the singlet scalar model. In fact we realize that regions with DM-nucleon cross section below the XENONnT limit and below the neutrino floor open up. In these regions the coupling $\kappa_{22}$ picks up small values while the coupling $\kappa_{12}$ and $\kappa_{11}$ can take large values. As can be seen in Fig. 2, Fig. 3 and Fig. 4, we have $\kappa_{12}\gg\kappa_{22}$ and $\kappa_{11}\gg\kappa_{22}$. This is plausible because now DM annihilation via $t$- and $u$-channel is kinematically allowed, and therefore a large value for $\kappa_{12}$ (and $\kappa_{11}$ in case of co-annihilation) regulates the relic density to its right value while the coupling $\kappa_{22}$ being a small value has a little impact on the relic density. Moreover, we emphasize that the co-annihilation processes are included in our computation for relic density, given that based on the standard lore these effects are effective when $\delta/m_{\text{DM}}\lesssim 10\%$.
Figure 2: The DM-nucleon cross section as a function of the DM mass is shown. Upper bounds from XENON1T and XENONnT, and lower bound from neutrino froor are placed. The ratio $\kappa_{11}/\kappa_{22}$ and ratio $\kappa_{12}/\kappa_{22}$ are shown as color spectrum in the left panel and right panel, respectively. Here $\alpha_{2}=0.25$.
Why are we encouraged to go beyond tree level in DM-nucleon scattering cross section? The standard expectation is that the tree level contribution is the dominant part in the cross section. According to the results presented in this section we find out that this is not the case for some regions in the parameter space. For points residing below the XENONnT bound and neutrino floor, we see that $\kappa_{12}\gg\kappa_{22}$ and $\kappa_{11}\gg\kappa_{22}$. Therefore, Feynman diagrams at loop level involving only the couplings $\kappa_{12}$ and $\kappa_{11}$, might have quite sizable contributions to the cross section in the regions with small DM-nucleon cross section. We will dub these loop effects as leading quantum corrections. As we will see in the next section these contributions are only achievable from triangle and box diagrams.

### v2

@@
We investigate the effect of quantum corrections on the elastic scattering cross section of dark matter off nucleus in two-scalar dark matter model. Among two extra singlet scalars in the two-scalar model, the lighter one is stable and plays the role of dark matter candidate and the heavier one contributes in dark matter co-annihilation processes in thermal history of the early universe. It is already known that the two-scalar model at tree level, unlike the single-scalar dark matter model, can easily evade the bounds from direct detection (DD) experiments. The claim here is that taking into account the loop effects, in some regions of the parameter space, the DM-nucleon cross section becomes larger than the tree level contribution. Therefore, loop effects move the regions which were below the neutrino floor at tree level, up to the regions which are detectable by future DD experiments.
@@ 1 Introduction
There are noticeably two classes of models with dark matter candidates escaping the strong bounds from DD experiments. One scenario is that the scattering cross section of DM-nucleon tends to zero at tree level due to some symmetry breaking structure, for example, a pseudo-Goldstone boson as DM candidate in the complex scalar model with softly broken symmetry [7, 8, 9]. As such, in models with scale symmetry breaking, the DD cross section at tree level may be reduced significantly at some regions in the parameter space [10, 11]. The second avenue deals with models wherein the DM-nucleon scattering cross section is momentum or velocity suppressed, giving rise to DM candidates evading DD bounds easily. Among models of this type we can recall thermal DM candidates which interact with nucleons through pseudoscalar operators [12, 13, 14, 15, 16, 17, 18, 19].
@@ 2 Two-Scalar Model
$\phi_{1}=\varphi_{1}\sin\epsilon+\varphi_{2}\cos\epsilon\,,~{}~{}~{}\phi_{2}=\varphi_{1}\cos\epsilon-\varphi_{2}\sin\epsilon\,,$ (4)
@@
We take the scalar field $\phi_{2}$ with smaller mass $m_{2}$ as our DM candidate. We note that the interaction terms $\lambda_{1}\varphi_{1}^{4}$, $\lambda_{2}\varphi_{2}^{4}$, $\lambda_{12}\varphi_{1}^{2}\varphi_{2}^{2}$, $\lambda_{13}\varphi_{1}\varphi_{2}^{3}$ and $\lambda_{31}\varphi_{1}^{3}\varphi_{2}$, do not contribute to the DM annihilation cross section as well as in the DM-nucleon cross section at tree level. However, in case we pick negative values for all the couplings $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{12}$, to be assured of vacuum stability, we may set the couplings of the above five interaction terms at fix values other than zero. Since our plan in this research is to include the leading quantum corrections we will see that these five interaction terms do not play a role. Therefore in this model we are left with five independent free parameters: $m_{1},m_{2},\alpha_{1},\alpha_{2}$ and $\alpha_{12}$. The scalar mass difference defined as $\delta=m_{1}-m_{2}>0$ is used in our computation. In the following sections we may use the identity $m_{2}\equiv m_{\text{DM}}$.
As a requirement for the vacuum stability, the potential part of the Lagrangian should fulfil the bounded from below condition. The relevant formulas are found in [41] for a generic potential with two scalars:
@@
Since in the present study we have set, $\lambda_{1},\lambda_{2},\lambda_{12},\lambda_{13}$ and $\lambda_{31}=0$, the parameters $a,b,c,d,e$ suited to our conditions read
$a=-\frac{\alpha_{1}^{2}}{4\lambda_{H}},~{}b=-\frac{\alpha_{1}\alpha_{12}}{\lambda_{H}},~{}c=-\frac{2\alpha_{12}^{2}+\alpha_{1}\alpha_{2}}{2\lambda_{H}},~{}d=-\frac{\alpha_{2}\alpha_{12}}{\lambda_{H}},~{}e=-\frac{\alpha_{2}^{2}}{4\lambda_{H}}.$ (8)
@@
where $\Gamma^{\text{SM}}_{higgs}$ is the Higgs decay width in the SM, $\Theta$ is the step function and $m_{h}\sim 125$ GeV is the Higgs mass. The decay width of the Higgs in three different channels are provided by
$\Gamma(h\to\phi_{1}\phi_{1})=\frac{(\alpha_{1}\sin^{2}\epsilon+\alpha_{2}\cos^{2}\epsilon+2\alpha_{12}\sin\epsilon\cos\epsilon)^{2}v^{2}}{8\pi m_{h}}\sqrt{1-4m_{1}^{2}/m_{h}^{2}}\,,$ (10)
$\Gamma(h\to\phi_{2}\phi_{2})=\frac{(\alpha_{1}\cos^{2}\epsilon+\alpha_{2}\sin^{2}\epsilon-2\alpha_{12}\sin\epsilon\cos\epsilon)^{2}v^{2}}{8\pi m_{h}}\sqrt{1-4m_{2}^{2}/m_{h}^{2}}\,,$ (11)
$\Gamma(h\to\phi_{1}\phi_{2})=\frac{((\alpha_{1}-\alpha_{2})\sin\epsilon\cos\epsilon+\alpha_{12}\cos 2\epsilon)^{2}v^{2}}{8\pi m_{h}^{3}}\sqrt{m_{h}^{2}-(m_{1}+m_{2})^{2}}\sqrt{m_{h}^{2}-(m_{1}-m_{2})^{2}}\,.$ (12)
The experimental upper limit at $95\%$ CL is found on the invisible Higgs decay, such that Br($h\to$ invisibles) $\lesssim 0.18$ [45]. This latter constraint becomes more effective for small mass of the scalars or large mixing angle.
It is also necessary to estimate the decay width of the heavier scalar. When the mass difference of the two scalars is smaller than the Higgs mass, $\delta<m_{h}$, then the decay $\phi_{1}\to\phi_{2}\bar{f}f$ is mediated by an off-shell Higgs particle. The SM fermion is indicated by $f$. The decay width is obtained as
[…]
where $t$ and $u$ are the Mandelstam variables in the relevant decay kinematics, and $N_{c}$ is the number of color for the SM fermion. The decay life-time of the heavy scalar is $\tau=\Gamma^{-1}$. We apply the code CalcHEP [46] to evaluate the decay width numerically. The scalar life-time for a set of parameters with reasonable magnitudes, $\delta=10$ GeV, $m_{1}=300$ GeV and $\alpha_{1},\alpha_{2},\alpha_{12}\sim 0.5$, is estimated as $\tau\sim 10^{-9}$ sec. It is also found that for larger mass of the heavy scalar, the life-time does not change significantly. Furthermore, we have verified that for larger values of $\delta$, the decay life-time becomes smaller because of the larger available phase space. The life-time of the heavy scalar is very smaller than the age of the universe in the regions of the parameter space which is relevant in this study.
@@ 3 Annihilation Cross Section and DM-nucleon cross section at Tree Level
The aim of this section is to reanalyze and update what is found within the present model in [37] from the calculations of the relic density and the DM-nucleon cross section at tree level. The DM density today depends on the so-called freeze-out temperature in the early universe. Around the freeze-out temperature the Hubble expansion rate exceeds the DM annihilation rate and on the other hand, the kinetic energy of the SM particles is low enough which leads to the suppression of the DM production. At this epoch the DM particles become non-relativisitc and go out of equilibrium and thus the DM density remains unchanged afterwards. The time evolution of the DM number density depends on the (co)annihilation processes of the two WIMPs. In the following we outline these (co)annihilation processes.
@@
There are three channels through which the scalar DM can annihilate. 1) Through $s$-channel by Higgs particle as a mediator; with the SM fermions, gauge bosons and the Higgs in the final state. 2) Through $t$- and $u$-channel by one of the scalars as mediator; with a pair of the Higgs particles in the final state. 3) Annihilation into a pair of Higgs particles via a contact interaction. In the left panel of Fig. 1, the corresponding Feynman diagrams for the DM annihilation are shown. We have not shown diagrams with more than two particles in the final state (i.e., $\phi_{2}\phi_{2}\to hhh$).
In addition, co-annihilation diagrams will be obtained by simply replacing one $\phi_{2}$ by one $\phi_{1}$ in the initial states. If we replace the two $\phi_{2}$ fields with $\phi_{1}$ fields, the annihilation diagrams for the heavier scalar will be obtained. In a model with two WIMPs, in principle one should solve two coupled Boltzmann equations to obtain the time evolution of the number density of each scalar. However, in practice it will be sufficient to solve a single Boltzmann equation (sum of the two coupled equations) by considering effective annihilation and co-annihilation cross sections [47, 48]. In the sum, terms describing the conversion processes and terms describing the decay processes cancel each other [48]. Therefore the sum of the number densities of the two scalars, $n=n_{1}+n_{2}$, will change via (co)annihilation of the two scalars, whose governing Boltzmann equation reads,
$\frac{dn}{dt}+3Hn=\langle\sigma_{\text{eff}}v\rangle(n_{eq}^{2}-n^{2})\,,$ (14)
where $\langle\sigma_{\text{eff}}v\rangle$ is the the thermal average over effective cross section times the relative velocity of DM particles at temperature $T$. The effective cross section is given by the following expression,
$\sigma_{\text{eff}}=g_{\text{eff}}^{-1}\Big{[}\sigma^{22}_{\text{ann}}+\sigma^{11}_{\text{ann}}(1+\frac{\delta}{m_{2}})^{3}e^{-2\delta/T}+2\sigma^{12}_{\text{co-ann}}(1+\frac{\delta}{m_{2}})^{3/2}e^{-\delta/T}\Big{]}\,,$ (15)
where $\sigma^{22}_{\text{ann}}$, $\sigma^{11}_{\text{ann}}$ and $\sigma^{12}_{\text{co-ann}}$, indicate the annihilation cross section of DM, annihilation cross section of heavy scalar and co-annihilation cross section, respectively. The effective degrees of freedom is, $g_{\text{eff}}=1+(1+\delta/m_{2})^{3/2}e^{-\delta/T}$.
@@
Let us assume for a moment that $\alpha_{12}=0$ (thus $\epsilon=0$). This will turn the model back into its simplest form, i.e. only one singlet scalar in the model. In this simple case the annihilation cross section contains terms each of them proportional to $\alpha_{2}^{2}$, $\alpha_{2}^{3}$ or $\alpha_{2}^{4}$. Thus the DM relic density is inversely proportional to these couplings. Now looking at the relevant vertex one can easily see that the DM-nucleon cross section depends on the same coupling, $\alpha_{2}$. In the right panel of Fig. 1 the diagram for the DM-nucleon elastic scattering is depicted. Since the DM-nucleon cross section and DM relic density dependency on the couplings go in opposite way, when varying the coupling $\alpha_{2}$, as it is shown in [37], it is not possible to respect both the DD bounds and the observed relic density at the same time (except for a resonance region where $m_{2}\sim m_{h}/2$).
@@
Now set $\alpha_{12}\neq 0$, and we get back to our present model. This brings in new interactions which give pivotal contribution to the DM annihilation cross section. We first identify three couplings related to the scalar-Higgs interactions. The strength of the interaction vertex $\phi_{2}\phi_{2}h$ is equal to $v\kappa_{22}$ with $\kappa_{22}=\alpha_{1}\cos^{2}\epsilon+\alpha_{2}\sin^{2}\epsilon-2\alpha_{12}\cos\epsilon\sin\epsilon$, the strength of the interaction vertex $\phi_{1}\phi_{2}h$ is equal to $v\kappa_{12}$ with $\kappa_{12}=\alpha_{12}\cos 2\epsilon+\cos\epsilon\sin\epsilon(\alpha_{1}-\alpha_{2})$, and finally the strength of the interaction vertex $\phi_{1}\phi_{1}h$ is $v\kappa_{11}$ with $\kappa_{11}=\alpha_{1}\sin^{2}\epsilon+\alpha_{2}\cos^{2}\epsilon+2\alpha_{12}\cos\epsilon\sin\epsilon$. Now the DM-nucleon cross section at tree level is only dependent on the coupling $\kappa_{22}$ while the DM annihilation cross section depends on the three couplings $\kappa_{11}$, $\kappa_{22}$ and $\kappa_{12}$. As a result, new viable regions in the parameter space might show up. The reason hinges on the fact that by incorporating these new contributions it becomes feasible to get small DM-nucleon cross section (small $\kappa_{22}$) respecting the current DD bounds, and in order to get the DM relic density right one can regulate the couplings $\kappa_{12}$ and $\kappa_{11}$ accordingly. In Appendix A, the annihilation cross sections are given for the full model. The DM-nucleon elastic scattering cross section at tree level is given in Appendix B.
@@
$0\leq\alpha_{1}\leq 1,~{}~{}~{}0\leq\alpha_{12}\leq 2,~{}~{}~{}10\leq\delta\leq 100~{}(\text{GeV}),~{}~{}~{}30\leq m_{2}\leq 2000~{}(\text{GeV})\,.$ (16)
We present our numerical results for the DM-nucleon cross section in terms of the DM mass for $\alpha_{2}=0.25$, $\alpha_{2}=1$ and $\alpha_{2}=2$ in Figs. 2-4, while all the constraints mentioned above are applied. In these plots some new features appear when DM mass exceeds $\sim 125$ GeV. These were absent in the singlet scalar model. In fact we realize that regions with DM-nucleon cross section below the XENONnT limit and below the neutrino floor open up. In these regions the coupling $\kappa_{22}$ picks up small values while the coupling $\kappa_{12}$ and $\kappa_{11}$ can take large values. As can be seen in Fig. 2, Fig. 3 and Fig. 4, we have $|\kappa_{12}/\kappa_{22}|\gg 1$ and $|\kappa_{11}/\kappa_{22}|\gg 1$. This is plausible because now DM annihilation via $t$- and $u$-channel is kinematically allowed, and therefore a large value for $\kappa_{12}$ (and $\kappa_{11}$ in case of co-annihilation) regulates the relic density to its right value while the coupling $\kappa_{22}$ being a small value has a little impact on the relic density. Moreover, we emphasize that the co-annihilation processes are included in our computation for relic density, given that based on the standard lore these effects are effective when $\delta/m_{\text{DM}}\lesssim 10\%$.
Figure 2: The DM-nucleon cross section as a function of the DM mass is shown. Upper bounds from XENON1T and XENONnT, and lower bound from neutrino floor are placed. The ratio $\kappa_{11}/\kappa_{22}$ and the ratio $\kappa_{12}/\kappa_{22}$ are shown as color spectrum in the left panel and right panel, respectively. Here $\alpha_{2}=0.25$.
Why are we encouraged to go beyond tree level in DM-nucleon scattering cross section? The standard expectation is that the tree level contribution is the dominant part in the cross section. According to the results presented in this section we find out that this is not the case for some regions in the parameter space. For points residing below the XENONnT bound and neutrino floor, we see that $|\kappa_{12}/\kappa_{22}|\gg 1$ and $|\kappa_{11}/\kappa_{22}|\gg 1$. Therefore, Feynman diagrams at loop level involving only the couplings $\kappa_{12}$ and $\kappa_{11}$, might have quite sizable contributions to the cross section in the regions with small DM-nucleon cross section. We will dub these loop effects as leading quantum corrections. As we will see in the next section these contributions are only achievable from triangle and box diagrams.
@@
Before we move on further, a comment is appropriate to mention. We may think that same type of quantum corrections is involved in the annihilation cross section. In fact what matters here is what we are comparing these corrections to. In the regions of interest (regions with very small DD cross section) we found that $|\kappa_{12}|\gg|\kappa_{22}|$. Looking at the DM annihilation diagrams at tree level we see two kind of diagrams: those involving the coupling $\kappa_{22}$ and those involving $\kappa_{12}$. In the regions of interest, the annihilation diagrams involving $\kappa_{12}$ plays the major role since $|\kappa_{12}|\gg|\kappa_{22}|$. Now, loop corrections will modify the tiny coupling $\kappa_{22}$ as $\kappa_{22}\to\kappa_{22}+\delta\kappa_{22}$, where $\delta\kappa_{22}$ comes from loop triangle diagrams involving $\kappa_{12}$. We note here that $\delta\kappa_{22}$ is much larger than $\kappa_{22}$ but quite smaller than $\kappa_{12}$. Even after this change and working with the effective coupling, the dominant annihilation diagrams are still those involving $\kappa_{12}$ because $\kappa_{12}$ is quite larger than $\kappa_{22}+\delta\kappa_{22}$. Therefore, the quantum corrections in the annihilation diagrams will not change meaningfully the picture we concluded based on our numerical computations. On the other hand, in the elastic scattering process we have a tree-level diagram involving $\kappa_{22}$. What we found indicates that the loop diagrams involving the effective coupling $\delta\kappa_{22}$ will enhance the DD cross section because $\delta\kappa_{22}$ can be quite larger than $\kappa_{22}$.
Moreover, non-zero values of the quartic couplings may add loop correction of order $\sim\lambda_{12}\kappa_{11}/(16\pi^{2})$ to the coupling $\kappa_{22}$ and thus it has effect on the DM-nucleus cross section. In this study we assumed that all the quartic couplings are negligible, i.e $\lambda_{i}\sim 0$, so the DM-nucleus cross section receives no contribution from this type of quantum correction.